# Galaxy formation with Wave/Fuzzy Dark Matter: The core-halo structure

### v1

@@ 5.2.4 Comparison with observations
We proceed by comparing the extracted profile with two known dwarf galaxies, Cetus and Andromeda XXI, that orbit the Milky Way and Andromeda, respectively, and exhibit the most similar density gaps (See Figure 7). In the left panel, the profiles are normalized based on their stellar peak density, while the right panel is additionally normalized by their respective transition radii. The black profiles represent the mean stellar profile of all Local Group dwarf spheroidal galaxies (Pozo et al., 2023). The right panel demonstrates that the extracted DM stellar profile (blue profile) aligns well with a real observed stellar profile, Andromeda XXI (red profile). This indicates that the extracted profile can reproduce observed dSph structures, in contrast to the CDM profiles.
@@
Furthermore, this panel reveals that the inner portions of the profiles are in good agreement, suggesting that they are consistent with a solitonic core. However, it’s crucial to note that Andromeda XXI and the extracted profile are at different redshifts, 5.5 and 0 respectively. This indicates that the extracted profile still has much more time to evolve, and its final core-halo structure at redshift 0 could substantially change over time.
| Halos | (-2<[Fe/H]<0) | (-4<[Fe/H]<-2) | (-6<[Fe/H]<-4) | (-8<[Fe/H]<-6) | |
| DM | 12340 | 93.12 | 4.91 | 0.71 | 1.26 |
| DM | 1950 | 80.62 | 14.82 | 1.44 | 3.12 |
| DM | 1550 | 77.22 | 16.91 | 1.68 | 4.19 |
| CDM | 21206 | 95.62 | 3.61 | 0.36 | 0.41 |
| CDM | 7572 | 88.55 | 9.17 | 1.12 | 1.16 |
| CDM | 3417 | 83.79 | 12.38 | 1.82 | 2.01 |
| WDM | 28328 | 89.48 | 8.43 | 0.76 | 1.33 |
| WDM | 3793 | 79.25 | 14.79 | 2.14 | 3.82 |
| WDM | 5122 | 78.21 | 16.34 | 1.44 | 4.01 |
| Halos | ||||||||
| DM | 12340 | 11.57 | 8.02 | 0.92 | 48.72 | 37.41 | 10.46 | 3.41 |
| DM | 1950 | 5.66 | 2.82 | 0.16 | 72.10 | 16.41 | 7.23 | 4.26 |
| DM | 1550 | 5.10 | 2.89 | 0.15 | 64.45 | 27.36 | 8.19 | 0 |
| CDM | 21206 | 13.21 | 5.58 | 0.74 | 48.60 | 37.27 | 7.75 | 6.38 |
| CDM | 7572 | 24.57 | 8.24 | 2.02 | 29.32 | 30.78 | 14.05 | 25.85 |
| CDM | 3417 | 7.99 | 10.87 | 0.87 | 35.91 | 24.00 | 9.51 | 30.58 |
| WDM | 28328 | 24.56 | 13.75 | 3.37 | 38.04 | 36.55 | 12.05 | 13.36 |
| WDM | 3793 | 10.18 | 6.14 | 0.63 | 63.88 | 20.43 | 6.52 | 9.17 |
| WDM | 5122 | 16.08 | 9.31 | 1.50 | 49.57 | 38.54 | 11.21 | 0.68 |
5.3 Stellar Morphology
In this section, we will compare the stellar distribution of each galaxy across different DM simulations. Figure 8 presents the stellar morphology at three projections for the isolated G1 at . The main panels in this plot display the stellar distribution for the entire galaxy at a scale of approximately 55 kpc/h, while the additional sub-panels provide a closer view of the central regions of the galaxy (5 kpc/h). This allows for a detailed comparison between different stellar cores.
Figure 8 distinctly illustrates variations among three DM models, revealing that stars exhibit a more asymmetric structure in DM compared to a slightly symmetric shape seen in WDM and progressively a more symmetric shape in CDM, aligning well with theoretical predictions (Gao & Theuns, 2007; Schive et al., 2020). Notably, there is a significant discernible difference between the stellar distributions in WDM and DM. Despite appearing similar in scale in Figure 1, stars in WDM are actually more flattened than in DM. Upon closer examination in the subpanels with zoomed data, stars exhibit a spherical form in the DM core, but a more elliptical shape towards the center for both CDM and WDM. The plot effectively highlights the disparities between the core and cuspy centers, indicating that stars tend to concentrate in smaller cores in DM model, contrasting the cusps observed in CDM/WDM. Additionally, aside from having fewer stars than CDM and WDM, DM has a smaller proportion of stars in the central regions (see Table 4). This suggests a more extended stellar profile around the soliton, in line with recent discoveries of halos around some studied dwarfs (Chiti et al., 2021; Collins et al., 2021). These extended halos are inherent to DM, composed of excited states above the ground state soliton, as evidenced by the NFW form, as predicted by the DM simulations (Schive et al., 2014a). This reflects the fundamentally non-relativistic nature of DM. Moving to Figure 11, we observe WDM/CDM data for z 2.23. The results are comparable for both models when compared to the data from redshift z 5.56, showing an expected symmetry.
The additional sub-panels in Figure 8 prominently display the predicted asymmetry for DM, a distinctive feature not observed in the symmetric distributions of stars in WDM and CDM models (Schive et al., 2020). This asymmetry in DM arises from the solitonic characteristic ‘jumping’ behaviour (soliton oscillations) on the scale of the de Broglie wavelength , or the core radius (), owing to its standing wave nature (Li et al., 2021; Chiang et al., 2021). The jumps occur over a timescale of years for a dSph with 0.3 kpc and 10 km/s, values closely matching those for this galaxy (as referred in Tables 1 and 2 and check out Pozo et al. (2023) for a deeper explanation of ). Notably, this timescale is equivalent to the stellar crossing time of the soliton, . In essence, this means that stars ‘sense’ the movement of the soliton and respond accordingly, resulting in the observed asymmetry (Li et al., 2021; Chiang et al., 2021).
5.3.1 Stellar age vs metallicity
@@
Metallicity plays a crucial role in unraveling the galaxy history. Here, we explore the relationship between the stellar metallicity and age. The corresponding data is displayed in the right columns of Figures 9 and 12. These figures illustrate the [Fe/H] versus stellar age profiles for G1 at the and for galaxies 1 and 3 at the . We chose to focus on these specific galaxies to emphasize the most pertinent data associated with instances where the potential presence of a core-halo structure has been identified.
@@
We observe that in DM and WDM, the stellar metallicity exhibits a smoother decline as the stellar age increases. On the other hand, CDM displays a more pronounced breaks and a more discrete age vs. [Fe/H] profile, suggesting that its associated stars may have diverse origins. This disparity could arise from a combination of a few different factors. In CDM, the stellar cycle is more extended, with the first stars forming around z30 and continuing up to 9.8 billion years (as seen in Figure 9). In contrast, in WDM/DM, star formation begins around z13 (Mocz et al., 2019, 2020). Additionally, DM structures form earlier in CDM compared to WDM/DM, a consequence of the small-scale cut-off in the power spectrum.
@@
Continuing with this idea, all three galaxies exhibit a peak of star formation around 9.6 billion years and 9.4 billion years, with the former being the predominant one for WDM/DM and the latter for CDM. Additionally, CDM is the only model with an additional small peak at 9.5 billion years. It is evident that all three models display a moderate to strong negative correlation between stellar age and metallicity. Similar trends and predictions are confirmed in Figure 12 at .
@@
The observations in Figures 9 and 12 suggest that metal-rich stars tend to be more concentrated in the central regions of the galaxies, while metal-poor stars are located farther out, which is consistent with previous studies (Dietz et al., 2020; Santucci & otehrs, 2020). Additionally, the majority of stars are concentrated within distances compatible with the core-halo structure (r < 2 kpc, marked with a vertical dashed line). Of particular interest is that the galaxies where a compatible core-halo structure has been identified (DM at and WDM at ) exhibit a radially-offset region of stellar-density enhancement, which is especially pronounced in WDM at (as seen in Figure 12). Notably, WDM is the only case where this concentration of stars extends beyond the theoretical limit of the core-halo structure (r > 2 kpc). This peculiarity may be attributed to the filamentary environment observed for this galaxy, an environment that cannot be reproduced in the CDM model (as shown in Figure 1). In a future work we will study this in more detail. In summary, galaxies with a higher concentration of stars extending along their halos, rather than being centrally concentrated, seem to be more likely to develop the core-halo structure.
5.3.2 Spatial distribution of metallicity and age
@@
Figures 10 and 13 investigate the correlation between the mean stellar metallicity and the stellar radii for galaxies from different DM models at both redshifts. The mean stellar age is depicted in the color bars.
@@
The analysis of the correlation between stellar age and the spatial distribution of metallicity (as shown in Figures 10 and 13) reveals distinct differences among the three DM models. One notable distinction is that in CDM and WDM stars are relatively more metal-rich ([Fe/H] -1.0) compared to DM ([Fe/H] -1.25). Furthermore, CDM and WDM exhibit a smoother, continuous decline in metallicity from the center to the outer halo, whereas DM displays a more discrete profile, with noticeable drops in metallicity around 0.3 kpc and 0.5 kpc (as observed in the right panel of the first row in Figure 10). Regarding the spatial distribution of stellar age, WDM exhibits older stars concentrated in the center, while the mean age decreases towards the outer parts of the halo. This trend is opposite to what is observed for CDM and DM. However, it is important to note that for DM, the mean age is lower compared to the mean age observed for CDM, which aligns with the predicted stellar history between both models (Mocz et al., 2019, 2020). Tables 3 and 4 provide additional data regarding the percentage of metallicity and age in the stellar profiles of all the galaxies. These findings further contribute to our understanding of how different DM models influence the stellar properties of galaxies.
@@ 6 Discussion and Conclusions
The comparison between the core-halo structures of WDM and DM simulations with the CDM predictions reveals intriguing distinctions. In the case of WDM and DM, the core-halo structure appears to naturally stem from their respective power spectrum small-scale cut-off, whereas for CDM this structure is not anticipated. This notion could be further examined by tracking the stars of each galaxy across different epochs and observing their initial positions to scrutinize in more detail whether the stars appearing in the outer extended halos were born within their halo or in the outer filamentary structures. This observation, combined with other disparities between CDM predictions and observations, has spurred the investigation of alternative DM candidates like WDM and DM. Moreover, the fact that the stellar profiles in each DM scenario adhere to the predicted behaviours, with a cuspy profile for WDM and CDM, and a flat core structure for DM, supports the notion that stars trace DM. Figure 4 vividly illustrates that the DM, stellar, and the gaseous density profiles (especially stellar and DM) differ significantly in each DM scenario, and that the resulting stellar profile varies depending on the DM profile (cuspy/core discrepancy and core-halo structure). This underscores the potential significance of studying the properties of stars to achieve a deeper comprehension of the characteristics of DM.
@@
Nonetheless, it is also true that in CDM and WDM both of the stellar and DM profiles appear to shift from a cuspy profile to cored one when decreasing the redshift (though never becoming as flat as DM predictions, which exhibit a better consistency with the observations (Sanchez Almeida et al., 2023)). This transformation from cuspy to cored profiles for CDM/WDM seems to align with the asserted impact of baryonic feedback in generating cores from originally cuspy halos. Feedback induces bursts of star formation and outflows, modifying the DM distribution and mitigating small-scale issues in CDM (Read & Gilmore, 2005; Amorisco et al., 2013; Chan et al., 2015), potentially resolving the cuspy-core problem. Furthermore, the simulations conducted by Mocz et al. (2019, 2020) were highly informative, demonstrating that baryonic feedback may not possess sufficient strength to convert cusps into cores for z>6, but it does open up possibilities for lower redshifts. This coincides with our findings, where the cusps observed for CDM/WDM at z5.56 seem to be alleviated around z2.23.
@@
Certainly, the detection of the core-halo structure in WDM simulations, despite the absence of a soliton, is a topic of discussion. It seems that the smoothness of the cut-off in the initial dark matter power spectrum, which is likely more pronounced for WDM, may play a role in the formation of the core-halo structure in the stellar profiles. This suggests that the soliton itself may not be the sole factor determining the presence of the core-halo structure. However, it is worth noting that the presence of a soliton can be important as the transition feature in the DM and star profiles is a result of the power spectrum cut-off. The presence of a soliton may have a more notable impact on the early formation of the core-halo structure in DM (at ) compared to WDM (at ).
@@
Indeed, the small-scale cut-off of the initial power spectrum appears to be the pivotal factor in determining the core-halo structure, as it leaves its mark on both the DM and the stellar profiles. The resemblance of the large-scale structures between the WDM and DM, despite their distinct physics on smaller scales, lends support to this notion. It is worth noting that the underlying cause of the cut-off differs for WDM and DM, with WDM exhibiting a thermal free-streaming cut-off, while DM featuring a quantum mechanical cut-off. However, this distinction does not seem to significantly impact larger scales where the cut-off length scale is imprinted on the profiles. The simulation data presented by Mocz et al. (2020) demonstrates that both of the WDM and DM yield core-halo stellar profiles, which appear to be at odds with CDM predictions. Further analysis is warranted to validate this finding and to probe the hypothesis that the core-halo structure arises from the small-scale cut-off of the initial power spectrum.
@@
On the other hand, the discrepancies found in the stellar morphology, age, and metallicity suggest the notion of a less concentrated and non-cuspy inner halo as a prerequisite for the observed core-halo structure in local dwarfs. This idea gains further support when we observe the higher concentration of stars in CDM/WDM models compared to DM, which exhibit a lower proportion of stars in their halo centers, indicating more extended stellar profiles. Furthermore, the spatial distribution of metallicity and age indicates that the core-halo structure emerges in galaxies where the greater concentration of stars is not extremely close to the center, a condition that appears only feasible in a filamentary environment, which is challenging to replicate for CDM. Finally, the solitonic stellar imprint is substantiated by the observed asymmetry in the core of G1 in the DM scenario, as a consequence of the previously analyzed solitonic random walk (Li et al., 2021; Chiang et al., 2021), where stars are thought to “sense” the movement of the soliton and respond accordingly, resulting in the observed asymmetry, in stark contrast to CDM/WDM outcomes. This further strengthens the idea of distinct stellar behaviors for each DM model. Quantifying this effect with the deepest JWST images of different dwarfs at various redshifts could yield valuable insights. Additionally, a more detailed analysis of this soliton ‘jumping’ phenomenon could be carried out with data from lower redshifts, an area we plan to explore in future studies by tracking the motion of stars around the soliton across different redshifts.
@@
We conclude that CDM is under significant tension with core-halo observations. Moreover, there are several compelling reasons to favor DM over WDM/CDM. Firstly, as highlighted previously, DM seems to align better with observations in terms of the presence and structure of core-halo profiles in galaxy halos, as opposed to CDM. Additionally, DM predicts a granularity in the lensed halos Amruth et al. (2023) that has not been observed in WDM simulations. Furthermore, the lack of detection of WDM particles in laboratory searches is a substantial drawback for this model. In contrast, DM posits light bosons that would necessitate a much larger laboratory scale of 1 kpc for detection, making it challenging to rule out this possibility based solely on laboratory experiments. Overall, the combination of better agreement with observations and the lack of definitive evidence against it makes DM a promising candidate for dark matter, and a more viable alternative to WDM/CDM.
Acknowledgements

### v2

@@ 5.3.4 Comparison with observations
We proceed by comparing the extracted DM and CDM G1 profiles with two known dwarf galaxies, Cetus and Andromeda XXI, that orbit the Milky Way and Andromeda, respectively, and exhibit the most similar density gaps (see Figure 7). In the left panel, the profiles are normalized based on their stellar peak density, while the right panel is additionally normalized by their respective transition radii. The black profiles represent the mean stellar profile of all Local Group dwarf spheroidal galaxies (Pozo et al. 2023). The right panel demonstrates that the extracted DM stellar profile (blue profile) aligns well with a real observed stellar profile, Andromeda XXI (orange profile). This indicates that the extracted profile can reproduce observed dSph structures, in contrast to the CDM profile, here represented in red.
@@
Furthermore, this panel reveals that the inner portions of the profiles are in good agreement, suggesting that they are consistent with a solitonic core. However, it is crucial to note that Andromeda XXI and the extracted profile are at different redshifts, 5.5 and 0 respectively. This indicates that the extracted profile still has much more time to evolve, and its final core-halo structure at redshift 0 could substantially change over time.
5.4 Stellar age vs. metallicity
@@
Metallicity plays a crucial role in unraveling the galaxy history. Here, we explore the relationship between the stellar metallicity and age. The corresponding data is displayed in the right columns of Figures 8 and 11. These figures illustrate the [Fe/H] versus stellar age profiles for G1 at the and for Galaxies 1 and 3 at the . We chose to focus on these specific Galaxies to emphasize the most pertinent data associated with instances where the potential presence of a core-halo structure has been identified.
@@
We observe that DM and “WDM” have pretty similar profiles with the metal-poor concentrations at similar ages, relaying a similar stellar history. On the other hand, CDM displays a more pronounced breaks(more discrete) age vs. [Fe/H] profile, suggesting that its associated stars may have diverse origins. This disparity could arise from a combination of a few different factors. In CDM, the stellar cycle is more extended, with the first stars forming around z30 ( stellar age 0.2) and continuing up to 1.2 billion years (as seen in Figure 8). In contrast, in “WDM”/DM, star formation begins around z13 ( stellar age 0.5) (Mocz et al. 2019, 2020). Additionally, DM structures form earlier in CDM compared to “WDM”/DM, a consequence of the small-scale cutoff in the power spectrum.
@@
Continuing with this idea, all three Galaxies exhibit a peak of star formation around 0.6 billion years and 1.0 billion years, with the former being the predominant one for “WDM”/DM and the latter for CDM. Additionally, CDM is the only model with an additional small peak at 0.85 billion years. As the current data does not cover more than 1 billion years, we cannot confirm the expected negative correlation between stellar age and metallicity. Similar data and predictions are shown in Figure 11 at .
@@
The observations in Figures 8 and 11 suggest that metal-rich stars tend to be more concentrated in the central regions of the Galaxies, while metal-poor stars are located farther out, which is consistent with previous studies (Dietz et al. 2020; Santucci & otehrs 2020). Additionally, the majority of stars are concentrated within distances compatible with the core-halo structure (r ¡ 2 kpc, marked with a vertical dashed line). Of particular interest is that the Galaxies where a compatible core-halo structure has been identified (DMG1 at and “WDM”G3 at ) exhibit a radially offset region of stellar-density enhancement, which is especially pronounced in “WDM”G3 at (as seen in Figure 11). Notably, “WDM”G3 is the only case where this concentration of stars extends beyond the theoretical limit of the core-halo structure (r ¿ 2 kpc). This peculiarity may be attributed to the filamentary environment observed for this Galaxy, an environment that cannot be reproduced in the CDM model (as shown in Figure 1). In a future work we will study this in more detail. In summary, galaxies with a higher concentration of stars extending along their halos, rather than being centrally concentrated, seem to be more likely to develop the core-halo structure.
@@
Figures 9 and 12 investigate the correlation between the mean stellar metallicity and the stellar radii for Galaxies from different DM models at both redshifts. The mean stellar age is depicted in the color bars.
@@
The analysis of the correlation between stellar age and the spatial distribution of metallicity (as shown in Figures 9 and 12) reveals distinct differences among the three DM models. One notable distinction is that in CDM and “WDM” stars are relatively more metal-rich ([Fe/H] -1.0) compared to DM ([Fe/H] -1.25). Furthermore, CDM and “WDM” exhibit a smoother, continuous decline in metallicity from the center to the outer halo, whereas DM displays a more discrete profile, with noticeable drops in metallicity around 0.3 kpc and 0.5 kpc (as observed in the right panel of the first row in Figure 9). Regarding the spatial distribution of stellar age, “WDM” exhibits older stars concentrated in the center, while the mean age decreases toward the outer parts of the halo. This trend is opposite to what is observed for CDM and DM. However, it is important to note that for DM, the mean age is lower compared to the mean age observed for CDM, which aligns with the predicted stellar history between both models (Mocz et al. 2019, 2020). Tables 3 and 4 provide additional data regarding the percentage of metallicity and age in the stellar profiles of all the Galaxies. These findings further contribute to our understanding of how different DM models influence the stellar properties of galaxies.
@@ 6 Discussion and conclusions
The comparison between the stellar core-halo structures (not the stellar core itself, but the observed inner core and outer halo structure in Figures 3 and 6, first detected in Pozo et al. (2024)) of DM simulations with the CDM predictions reveals intriguing distinctions. In the case of DM, the core-halo structure appears to naturally stem from the also observable DM core-halo structure (see the first panel of the top row of figure 5), whereas for CDM, this structure is not anticipated. Moreover, as seen in figure 5, it is clear that in DM the N-body simulations exhibit a cuspy DM profile, whereas only DM presents solitonic cores. Any core formation in CDM and “WDM” seems to be driven by baryonic physics rather than being directly related to the DM component. Therefore, baryonic physics is the most likely explanation for the transformation of the observed DM cusp in “WDM”–CDM into stellar cores. This notion could be further examined by tracking the stars of each galaxy across different epochs and observing their initial positions to scrutinize in greater detail whether the stars appearing in the outer extended halos were born within their halo or in the outer filamentary structures. This observation, combined with other disparities between CDM predictions and observations, has spurred the investigation of alternative DM candidates such as “WDM” and DM. Moreover, the fact that the stellar profiles in each DM scenario follow the predicted behaviors, with a cuspy profile for CDM and a flat or prominent core structure for DM, supports the notion that stars could trace DM. Figure 5 vividly illustrates significant differences in the density profiles of DM, stars, and gas in each DM scenario. The resulting stellar profile varies depending on the DM model, indicating a cuspy–core discrepancy and core-halo structure. Finally, this work aims to clarify how the core-halo structure can evolve and how definitive it can be to support any dark matter model. It can be observed that the core-halo structure appears in the DM dark matter density profiles but never for CDM–“WDM”. This point alone might not be significant enough if it were not for the fact that such core-halo structures, which we are observing for dark matter, also appear in the DM Galaxy 1 stellar profile and probably in Galaxy 3 for “WDM” at z=2.23, as can be seen in Figures 3 and 4. However, it never appears for CDM. This strongly supports the assertion that the stellar profiles show discrepancies for the three different dark matter models, as we claim in this work.
@@
This underscores the potential significance of studying the properties of stars to achieve a deeper comprehension of the characteristics of DM. Nonetheless, it is also true that in CDM and “WDM” the stellar and DM profiles both appear to shift from a cuspy profile to a cored one when decreasing the redshift, although they never become as flat and prominent as the DM predictions, which exhibit a better consistency with the observations (Sanchez Almeida et al. 2023). This transformation from cuspy to cored profiles for CDM–“WDM” seems to align with the asserted impact of baryonic feedback in generating cores from originally cuspy halos. While in the case of “WDM”, the inner cores are claimed as well (Macciò et al. 2012; Lovell et al. 2014; Bode et al. 2001; Destri et al. 2013). Feedback induces bursts of star formation and outflows, modifying the DM distribution and mitigating small-scale issues in CDM (Read & Gilmore 2005; Amorisco et al. 2013; Chan et al. 2015), potentially resolving the cuspy-core problem. Furthermore, the simulations conducted by Mocz et al. (2019, 2020) were highly informative, demonstrating that baryonic feedback may not possess sufficient strength to convert cusps into cores for z¿6, but it does open up possibilities for lower redshifts. This coincides with our findings where the cusps observed for CDM–“WDM” at z5.56 seem to be alleviated around z2.23.
@@
Certainly, the possible detection of the stellar core-halo structure in “WDM” simulations, despite the absence of a soliton, is a topic of discussion. It seems that the smoothness of the cutoff in the initial dark matter power spectrum, which is likely more pronounced for “WDM”, may play a role in the formation of the core-halo structure in the stellar profiles. This suggests that the soliton-interference pattern change of regime itself may not be the sole factor determining the presence of the core-halo structure. Nevertheless, it is extremely important to note that this conclusion is a product of our poor understanding of why the stellar core-halo structure appears for “WDM” (at z = 2.23, Figure 4 Galaxy 3) as we thought that it could only appear as a consequence of the expected profile observed in all ideal DM dark matter models, featuring an inner flat core fitted with a soliton and an outer characteristic asymptotic profile. Even more analysis is thus required for this topic, and we conclude that it has to be a coincidence due to the absence of such structure for “WDM” dark matter density profiles (see figure 5).
@@
It is worth noting that the underlying cause of the cutoff differs for “WDM” and DM; “WDM” exhibits a thermal free-streaming cutoff, while DM features a quantum mechanical cutoff. However, this distinction does not seem to significantly impact larger scales where the cutoff length scale is imprinted on the profiles. The simulation data presented by Mocz et al. (2020) demonstrates that only the DM core-halo stellar profiles are reliable, due to the presence of this structure in both gas and DM, which is only noticeable in DM simulations.
@@
On the other hand, the discrepancies found in the stellar morphology, age, and metallicity suggest the notion of a less concentrated and noncuspy inner halo as a prerequisite for the observed core-halo structure in local dwarfs. This idea gains further support when we observe the higher concentration of stars in CDM–“WDM” models compared to DM, which exhibit a lower proportion of stars in their halo centers, indicating that the stellar profiles are more extended. Furthermore, the spatial distribution of metallicity and age indicates that the core-halo structure emerges in galaxies where the greater concentration of stars is not extremely close to the center, a condition that appears only feasible in a filamentary environment, which is challenging to replicate for CDM. Finally, the solitonic stellar imprint is substantiated by the observed asymmetry in the core of G1 in the DM scenario, as a consequence of the previously analyzed solitonic random walk (Li et al. 2021; Chiang et al. 2021), where stars are thought to sense the movement of the soliton and respond accordingly, resulting in the observed asymmetry, in stark contrast to CDM–“WDM” outcomes. This further strengthens the idea of distinct stellar behaviors for each DM model. Quantifying this effect with the deepest JWST images of different dwarfs at various redshifts could yield valuable insights. Additionally, a more detailed analysis of this soliton jumping phenomenon could be carried out with data from lower redshifts, an area we plan to explore in future studies by tracking the motion of stars around the soliton across different redshifts.
The light boson used here, 2.5 eV, is consistent with the canonical value based the internal properties of the classical dwarf spheroidal galaxies (Schive et al. 2014a; Chen et al. 2017; Pozo et al. 2023), which has been claimed to be significantly lighter than some independent constraints using the Ly-forest of eV (Iršič et al. 2017), but it should be noted that the forest estimates are not direct and rely on analogy with “WDM” simulations as they have similar power spectrum suppressions at high frequency (from the de Broglie scale and the free streaming scale, respectively); however, the DM interference introduces additional power on small scales that is pervasive, including the low-density regions probed by the Ly-forest, where a wide range of the de Broglie scale is seen in the simulations extending to low momentum in the voids (see Figure S2 of Schive et al. (2014a)). Furthermore, the claimed discrepancy with DM is below lines widths of and only at z5, in the steeply declining Jeans turnover where the gas temperature is degenerates with the DM power spectrum, and where spatial temperature variations from AGN and cluster galaxy gas heating sources (Doughty et al. 2025) adds to the forest variance, with widespread metal injection and wide forest gaps (Zhu. et al. 2022), thus raising the observed power spectrum; however, all this is too uncertain to confidently model in forest simulations. Hence, reliance on Ly-forest regarding the boson mass of Wave DM is far less direct than our lensing-based estimate (Amruth et al. 2023; Broadhurst et al. 2024). However, insisting on smaller scale power can be simply achieved with an additional heavier boson species (Rogers & Peiris 2021), as in the generic string axiverse (Arvanitaki et al. 2010). An additional heaver boson is also motivated by the compact scale of the new UFG class of dwarf galaxy where a boson mass for which a boson mass of eV is indicated dynamically (Pozo et al. 2023), forming a minority of the DM, most of which is comprised of the canonical eV boson species. Galaxy formation in this two-boson model has been shown to be viable, capable of forming UFG dwarfs (Nhan Luu et al. 2024).
A lensing based estimate of eV is claimed by Powell et al. (2023), who argue that their reconstructed source is fully resolved with VLBI, so that sufficiently light DM would cause disruption. The breadth of their predicted band of DM critical curves is wide, about 25% of the Einstein radius, whereas a band narrower by a factor of 3 is predicted by Eq. 3, and by comparable lens simulations of DM (Chan et al. 2020b; Amruth et al. 2023), which would leave the source unscathed, and hence insensitive to the lighter boson mass we estimate here.
A tight constraint of ¿3eV has been claimed dynamically by Dalal & Kravtsov (2022) using two dwarf galaxies Segue I and II. This claim rests mainly on Segue II; the significance of their boson mass estimate relies mainly on the viability of Segue II, because its velocity dispersion is too small to be detected, unlike the 5.4 Km/s dispersion established for Segue I. This reliance on Segue II is evident in Figure 4 of Dalal & Kravtsov (2022). However, this dwarf Segue II is widely recognized as an extreme tidal remnant in the careful work of Kirby et al. 2013, and follow-up work where its heavy progenitor origin is reinforced by star orbits(Simon 2019). Kirby (Kirby et al. 2013) point out that Segue II is an extreme outlier in terms of its high metallicity and low (undetected) velocity dispersion, which they conclude implies a much more massive origin similar to the regular dSph galaxy, and also conclude that Segue II may be the barest remnant of a tidally stripped Ursa Minor-sized galaxy. If this is the case, it is the best example of an ultra-faint dwarf galaxy that became ultra-faint through tidal stripping. Thus, it is widely believed that Segue II is not representative of the class of low-mass galaxies that Dalal & Kravtsov (2022) would like to use to constrain DM. On the other hand, Segue I appears to be a bona fide representative of the ultra-faint and dark matter-dominated dwarf with a well-detected velocity dispersion. A Jeans analysis of this dwarf finds a boson mass of eV (Pozo et al. 2023). This is an order of magnitude difference, and the UFD class as a whole indeed fits well with this higher boson mass. A case for two boson species has been made to encompass both UFD and dSPh classes by some of us (Pozo et al. 2023). The argument of Dalal & Kravtsov (2022) rests on simulations that seem to underpredict the observed velocity dispersion of 3.5 km/s, so that it is not clear if Segue II alone challenges DM. Instead, independent Jeans analysis shows that it is consistent in terms of boson mass with the other UFD galaxies with eV (Pozo et al. 2023) and distinct from the eV dominant light boson indicated by the dark matter dominated dSph class (Pozo et al. 2023). Two such light boson species are arguably well motivated by the axiverse scenario where a discrete mass spectrum of light bosons is generic to string theory Arvanitaki et al. (2010), and is the subject of recent multi-boson species simulations Nhan Luu et al. (2024); Chan et al. (2020b).
@@
We conclude that CDM is under relative tension with stellar core-halo observations (Pozo et al. 2024, 2023), as it is never detected for any simulated CDM halo. More analysis in different simulations needs to be done in the future to truly support the idea that CDM does not produce these stellar core-halo structures in galaxies. Moreover, there are several compelling reasons to favor DM over “WDM”–CDM. First, as highlighted previously, DM seems to align better with observations in terms of the presence and structure of core-halo profiles in galaxy halos, as opposed to CDM. Additionally, DM predicts a granularity in the lensed halos Amruth et al. (2023) that has not been observed in “WDM” simulations. Furthermore, the lack of detection of “WDM” particles in laboratory searches is a substantial drawback for this model. In contrast, DM posits light bosons that would necessitate a much larger laboratory scale of 1 kpc for detection, making it challenging to rule out this possibility based solely on laboratory experiments. Overall, the combination of better agreement with observations and the lack of definitive evidence against it makes DM a promising candidate for dark matter, and a more viable alternative to “WDM”–CDM.
Acknowledgements.